OpenStudy (anonymous):

What's the general form for the multiple angle.... cos x = -1/2 cos x = 0

OpenStudy (anonymous):

Use the unit circle, so so look for a value the x is -1/2, so 2pi/3, 4pi/3 look for value x is o, so pi/2,3pi/2 so final form use the period of cos 2pi x= 2pi/3, 4pi/3+n2pi x=pi/2+3pi/2+n2pi

OpenStudy (anonymous):

Thank you. Do you mind helping with 1/2 more?

OpenStudy (anonymous):

yes no problem, do you understand the problem?

OpenStudy (anonymous):

whts the 2nd problem

OpenStudy (anonymous):

I guess I'll tell you what I think and we'll see if I understand them. It's x = 2pi/3 , 5pi/6, 5pi/3, and 11pi/6 The interval is [0,2pi) Let me see what the general forms I think would be, one sec.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Remember use unit circle and the trig functions period

OpenStudy (anonymous):

Okay, no, I tried to follow what you said but I'm confused. Do you mind rephrasing it or something? :/ Sorry for the inconvenience, I usually pick these things up quickly.

OpenStudy (anonymous):

Np, I am happy to help, I just finished studying, so I will make sure you understand it The values you are given are radians from the unit circle, you need to find the values for sin and cos cos is x coordinate and sin is y coordinate so (x,y) is (cos,sin) example 2pi/3, look on unit circle and coordinates are (-1/2,rad3/2) and sin and cosine period is at 2pi 1.) cos x= -1/2+n2pi, sin x= rad3/2 + n2pi Let me know if you understand

OpenStudy (anonymous):

Sorry I am a slow typer

OpenStudy (anonymous):

No worries and I think I get it now but let's say there was an "interval", would that affect it in any way? I mean, would both of the general forms still be used or just one?

OpenStudy (anonymous):

No the interval is telling you the period of cos and sin 2pi all trig functions general period is 2pi, except for tan and cotangent which period is pi when you need to worry about the intervals is when you start doing more trig identited sin(u+v) so for this kind of problem you do not need to worry about it, just keep the general periods

OpenStudy (anonymous):

So if my teacher were to ask for the general form of the angle, I'd have to put both of the general forms right?

OpenStudy (anonymous):

I see here that it's: 5pi/6 = -sqrt3/2, 1/2 5pi/3 = 1/2, -sqrt3/2 11pi/6 = sqrt3/2, -1/2 As one example, would the 5pi/6 one be... cos x = -sqrt3/2+n2pi, sin sin = rad1/2 + n2pi ?

OpenStudy (anonymous):

yes exactly, good job!

OpenStudy (anonymous):

In the first problem you have the find all values of (cos,sin), if given coordinates first

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

in the first problem it gave us cos=-1/2 and cos=0 so in the unit circle we find when cosine is -1/2 and 0 for all the radian values which was 2pi/3, 4pi/3 for -1/2, and pi/2, 3pi/2 for 0, and we put it in the general formula In the seconf problem they gave us the radians such as 5pi/6 so we found the (cos,sin) (xy) coordinates (-rad3/2,1/2) and put it in the general form

OpenStudy (anonymous):

So you would basically be doing the same thing or is there something different?

OpenStudy (anonymous):

You would still plug the xy values in x=values+n2pi for both problems 1st problem it gave us cos=-1/2, cos=0 1.) use unit circle to find when cos is -1/2 and 0 2.) cos at -1/2, radian values 2pi/3, 4pi/3 3.) cos at 0, radian values pi/2, 3pi/2 4.) x= 2pi/3, 4pi/3 + n2pi for 1/2 coordinate 5.) x= pi/2, 3pi/2 + n2pi for 0 coordinate 6.) Remember cos is x coordinate (cos, sin) sin is y coordinate 2nd Problem we are given 2pi/3, 5pi/6, 5pi/3, 11pi/6 1.) It is saying when Cos radians is 2pi/3, 5pi/6, 5pi/3 11pi/6 and if we are finding Sin radians 2pi/3 5p/6, 5p/3 11pi/6 2.) Unit circle when Sin,Cos values at 2pi/3, 5pi/6, 5pi/3, 11pi/6 2pi/3 so Cos= -1/2, Sin=rad3/2, (-1/2, rad3/2) 5pi/6 so Cos= -rad3/2, Sin= 1/2 (-rad3/2,1/2) 5pi/3 so Cos= 1/2, Sin=-rad3/2 (1/2, -rad3/2) 11pi/6 so Cos= rad3/2, Sin= -1/2 (rad3/2, -1/2) 3.) Now plug in the values to the general form Cos at 2pi/3, x= -1/2 + n2pi, Sin y=rad3/2 +n2pi Cos at 5pi/6, x= -rad3/2 +n2pi, Sin y= 1/2 +n2pi Cos at 5pi/3, x= 1/2 +n2pi, Sin y= -rad3/2 + n2pi Cos at 11pi/6 x= rad3/2 +n2pi, Sin y= -1/2 + n2pi Let me know if this makes sense

OpenStudy (anonymous):

Perfect sense. Thank you. I hate to feel like I'm bothering so if you'd like, just help me with this last one and that's it or I'll just keep looking online for help. tan (3x) = 0 tan (x)-1 = 0

OpenStudy (anonymous):

Np I will help you

OpenStudy (anonymous):

kk 1 sec

OpenStudy (anonymous):

you want the coordinates for this one also, right?

OpenStudy (anonymous):

Just the forms if you'd like, since I can take the coordinate from them myself, right?

OpenStudy (anonymous):

kk 1 sec, yes let me make sure

OpenStudy (anonymous):

We are using the trig identites, right

OpenStudy (anonymous):

kk got it

OpenStudy (anonymous):

So tan(3x)=0, tan(x)-1=0 1.) tan(3x)=0 0/3 is 0 x= 0 2.) tan(x)-1= 0 tan(x)= 1 so sin/cos = tan Values that get you 1 for tan x= pi/4, 5pi/4 sin rad2/2/cos 2rad/2=1 sin -rad2/2/cos 2rad 2 = 1

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

yes i believe so. where is says "rad" i wouldnt actually include that right?

OpenStudy (anonymous):

i just wrrite the number? youve got your medal btw :)

OpenStudy (anonymous):

thanks!, rad is square root sign, I would keep it in the work, I am working on your other problem, if you need anymore help, just notify me and I will finish your other problem

OpenStudy (anonymous):

thanks a lot, love.

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

kk i got it